# thanks: These two authors contributed equallythanks: These two authors contributed equally

### v2

@@
Teresa W. Lo    Han Kyou James Choi    Dean Huang Department of Physics, University of Washington, Seattle, Washington 98195, USA    Paul A. Wiggins pwiggins@uw.edu Department of Physics, University of Washington, Seattle, Washington 98195, USA Department of Bioengineering, University of Washington, Seattle, Washington 98195, USA Department of Microbiology, University of Washington, Seattle, Washington 98195, USA
@@ Appendix A Detailed development of the RLTO model
In this section, we provide a detailed development of the RLTO model. First, we provide a derivation of the growth rate as a function of the model parameters (Sec. A.1) as well as other methods (Secs. A.3, A.11, A.14, and A.17). For each of the results discussed in the main paper, we provide more detailed analyses, which include both supplemental results (Secs. A.7, A.8, A.10, A.12, and A.15) that support the story described in the main paper, as well as supplemental discussions (Secs. A.2, A.4, A.5, A.6, A.13, and A.16).
A.1 Methods: The derivation of the RLTO growth rate
A.1.1 Metabolic load
In the absence of cell cycle arrest, we will assume that the cell cycle duration $\tau$ is proportional to the overall metabolic load of the messages, proteins, and other cellular components [1]. Focusing on a particular gene:
$\frac{\tau}{\delta\tau}=N_{0}=N_{0}^{\prime}+\lambda\mu_{m}+\mu_{p},$ (1)
@@ A.1.1 Metabolic load
where $\delta\tau$ is a constant with units of time, $N_{0}$ is the total metabolic load (in protein equivalents), $N_{0}^{\prime}$ is the total metabolic load of everything but gene $i$ (in units of protein equivalents), and $\lambda$ is the message cost, the total metabolic load associated with an mRNA molecule relative to a single protein molecule of the gene product. We will assume the parameters $N_{0}$, $\delta\tau$ and $\lambda$ are global (i.e. gene independent), while the message number $\mu_{m}$ and translation efficiency $\varepsilon$, and therefore the protein number:
$\mu_{p}=\varepsilon\mu_{m},$ (2)
are gene specific. The total metabolic load in protein equivalents is:
$N_{0}\equiv L_{0}+\sum_{i}\{(\lambda+\varepsilon)\mu_{m}\}_{i},$ (3)
where $L_{0}$ is the load of non-protein and message cellular contributions and the sum runs over all genes. The metabolic load of transcription for gene $i$ is $\lambda\mu_{m,i}$ and for translation, $\mu_{p,i}=\varepsilon_{i}\mu_{m,i}$.
@@
Although the global parameters $N_{0}$, $\delta\tau$ and $\lambda$ provide an intuitive representation of the model, the relative growth rate depends on fewer parameters. Let $k$ and $k_{0}$ be the growth rates in the presence and absence of the metabolic load of gene $i$. The relative growth rate is:
$\frac{k_{0}}{k}=1+(\Lambda+E)\mu_{m},$ (4)
@@
where we have introduced two new reduced parameters: the relative load, defined as $\Lambda\equiv\lambda/N_{0}$, represents the ratio of the metabolic load of a single message to the total load and the relative translation efficiency, defined $E\equiv\varepsilon/N_{0}$, which is the ratio of the number of proteins translated per message to the total metabolic load $N_{0}$. (Note that due to the high multiplicity $N_{0}\gg(\lambda+\varepsilon)\mu_{m}$, we can ignore the distinction between $N_{0}$ and $N_{0}^{\prime}$ in the demoninator.) If we neglect the difference between the total metabolic load and the number of proteins, the proteome fraction for gene $i$ is:
$\Phi_{p}=E\mu_{m}.$ (5)
@@
Both reduced parameters, $\Lambda$ and $E$ are extremely small. In E. coli, we estimate that both $\Lambda$ and $E$ are roughly $10^{-5}$ and they are smaller still for eukaryotic cells. (See Sec. A.17.)
A.1.2 Growth rate with stochastic arrest
@@
is the exponential mean where $T$ is the stochastic cell cycle duration [3, 2].
Let $P_{+}$ be the probability of growth. When the cells are growing, the cell cycle duration is given by the metabolic load cell-cycle duration $\tau$ (Eq. 1). Evaluating the expectation gives:
$\ln 2=-\ln P_{+}+k\tau,$ (8)
which can be rearranged to give an expression for the population growth rate $k$:
$k=\tau^{-1}\ln 2P_{+}.$ (9)
As expected, the growth rate goes down as the probability of growth $P_{+}$ decreases, stopping completely at $P_{+}=\frac{1}{2}$.
Figure S1: Four perspectives on the fitness landscape. In each landscape, the red circle represents the fitness optimum. The red dotted line represents the mean protein number equal to the protein threshold $n_{p}=10^{2}$. Here fitness is quantified by the log growth rate: $s=\ln k_{\rm max}/k$. Panel A: Mean protein number versus noise. Panel B: Mean protein number versus translation efficiency. Panel C: Mean message versus protein numbers. Panel D: Message number versus translation efficiency.
A.1.3 RLTO growth rate
@@
$\ln P_{+}$ $=$ $\sum_{i\in{\cal E}}\ln\left(1-{\rm Pr}\{N_{p}<n_{p}\}_{i}\right),$ (11)
$\approx$ $-\sum_{i\in{\cal E}}{\rm Pr}\{N_{p}<n_{p}\}_{i},$ (12)
$=$ $-\sum_{i\in{\cal E}}\textstyle\left[\gamma(\frac{\mu_{m}}{\ln 2},\frac{n_{p}}{\varepsilon\ln 2})\right]_{i},$ (13)
where $\gamma$ is the regularized incomplete gamma function and the CDF of the gamma distribution (see  D.1.2 and D.1.4). Finally, we will be interested in the quantity:
$\ln\ln 2P_{+}\approx\ln\ln 2+-\sum_{i\in{\cal E}}\textstyle\frac{1}{\ln 2}\left[\gamma(\frac{\mu_{m}}{\ln 2},\frac{n_{p}}{\varepsilon\ln 2})\right]_{i},$ (14)
where we have again expanded the natural logarithm in the limit that the CDF is small.
A.1.4 When the ln approximation is avoided…
The approximation discussed in the previous section is extremely well justified at the optimal central dogma parameters; however, there are a set of figures where we cannot use it. In the fitness landscape figures (Fig. S1), we compute the fitness not just at the optimal values but far from them. Here we cannot approximate the natural log and we use the full expression in Eq. 11.
A.1.5 Single-gene equation
By combining Eqs. 1, 9, and 14, we can write an expression for the growth rate:
$\ln k$ $=$ $-\ln\left(L_{0}+\sum_{i}\left[(\lambda+\varepsilon)\mu_{m}\right]_{i}\right)+...$ (15)
$+\ln\ln 2-\sum_{i\in{\cal E}}\textstyle\frac{1}{\ln 2}\left[\gamma(\frac{\mu_{m}}{\ln 2},\frac{n_{p}}{\varepsilon\ln 2})\right]_{i}.$
We will work in the large multiplicity limit where each of the arrest probabilities is very small. To optimize the expression of gene $i$ we can drop all constant terms and write the simplified expression:
$\ln k$ $=$ $-\ln\left(N_{0}^{\prime}+(\lambda+\varepsilon)\mu_{m}\right)+...$ (16)
$+\ln\ln 2-\textstyle\frac{1}{\ln 2}\gamma(\frac{\mu_{m}}{\ln 2},\frac{n_{p}}{\varepsilon\ln 2})+...,$
where we have made the essential gene $i$ subscript implicit and we consider a total load minus gene $i$:
$N_{0}^{\prime}\equiv L_{0}+\sum_{j\neq i}\left[(\lambda+\varepsilon)\mu_{m}\right]_{j},$ (17)
interpreted as the metabolic load of all genes but $i$. However, we will work in the large multiplicity limit where we ignore the distinction between $N_{0}$ and $N_{0}^{\prime}$. The fitness landscape for different gene expression parameters is shown from four different perspectives in Fig. S1.
A.1.6 Summary of RLTO parameter values for figures
The parameter values for the RLTO model used for each figure in the paper are shown in Tab. S1. All calculations are independent of $\delta\tau$ and $L_{0}$.

### v3

@@ Appendix A Detailed development of the RLTO model
In this section, we provide a detailed development of the RLTO model. First, we describe the stochastic kinetic model for the central dogma, which introduces key quantities for the RLTO model (Sec. A.1). Next, we provide a derivation of the growth rate as a function of the model parameters (Sec. A.2) as well as other methods (Secs. A.4, A.12, A.15, and A.18). For each of the results discussed in the main paper, we provide more detailed analyses, which include both supplemental results (Secs. A.8, A.9, A.11, A.13, and A.16) that support the story described in the main paper, as well as supplemental discussions (Secs. A.3, A.5, A.6, A.7, A.14, and A.17).
A.1 Methods: Detailed description of the noise model
A.1.1 Stochastic kinetic model for the central dogma.
The equilibrium Paulsson model for the central dogma describes multiple steps in the gene expression process [12, 13, 9]: Transcription generates mRNA messages. These messages are then translated to synthesize the protein gene products [11]. Both mRNA and protein are subject to degradation and dilution [30]. At the single cell level, each of these processes are stochastic. We will model these processes with the stochastic kinetic scheme [11]:
[…]
where $\beta_{m}$ is the transcription rate (s-1), $\beta_{p}$ is the translation rate (s-1), $\gamma_{m}$ is the message degradation rate (s-1), and $\gamma_{p}$ is the protein effective degradation rate (s-1). The message lifetime is $T_{m}\equiv\gamma_{m}^{-1}$. For most proteins in the context of rapid growth, dilution is the dominant mechanism of protein depletion and therefore $\gamma_{p}$ is approximately the growth rate [31, 32, 9]: $\gamma_{p}=T^{-1}\ln 2$, where $T$ is the doubling time.
A.1.2 Statistical model for protein abundance.
To study the stochastic dynamics of gene expression, we used a stochastic Gillespie simulation [33, 34]. (See Sec. A.1.3.) In particular, we were interested in the explicit relation between the kinetic parameters $(\beta_{m},\gamma_{m},\beta_{p},\gamma_{p})$ and experimental observables.
Consistent with previous reports [12, 13], we find that the distribution of protein number per cell (at cell birth) was described by a gamma distribution: $N_{p}\sim\Gamma(a,\theta)$, where $N_{p}$ is the protein number at cell birth and $\Gamma$ is the gamma distribution which is parameterized by a scale parameter $\theta$ and a shape parameter $a$. (See Sec. A.1.4.) We refer to this distribution as the Paulsson model; The relation between the four kinetic parameters and these two statistical parameters has already been reported, and have clear biological interpretations [13]: The scale parameter:
$\theta=\varepsilon\ln 2,$ (10)
is proportional to the translation efficiency:
$\varepsilon\equiv\textstyle\frac{\beta_{p}}{\gamma_{m}},$ (11)
$a=\textstyle\frac{\beta_{m}}{\gamma_{p}};$ (12)
$\mu_{m}\equiv\beta_{m}T=a\ln 2,$ (13)
A.1.3 Gillespie simulation of stochastic kinetic scheme
Protein distributions based on the kinetic scheme defined in Sec. A.1.1 were simulated with a Gillespie algorithm, with specific parameter values for E. coli. Assuming the lifetime of the cell cycle ($T=30$ min) [35], mRNA lifetime ($T_{m}=2.5$ min) [36], and translation rate ($\beta_{p}\approx 500$ hr-1), the protein distributions for several mean expression levels were numerically generated for exponential growth with 100,000 stochastic cell divisions, with protein partitioned at division following the binomial distribution.
The gamma distributions for each mean message number with scale and shape parameters determined by the corresponding translation efficiency and message number ($\theta=\varepsilon\ln 2$, $a=\frac{\mu_{m}}{\ln 2}$) as used for the Gillespie simulation were also plotted with the protein distributions. We observe an excellent match between these Gillespie simulations and the Paulsson statistical noise model (i.e. gamma function) as shown in Fig. S1.
A.1.4 Gamma function and distribution conventions
$X\sim\Gamma(a,\theta),$ (14)
where $\Gamma(a)$ is the gamma function. The CDF is therefore:
$=$ $P_{\Gamma}(\textstyle\frac{x}{\theta}|a,1),$ (17)
$=$ $\gamma(a,x/\theta),$ (19)
where $\gamma$ is the regularized incomplete gamma function.
A.2 Methods: The derivation of the RLTO growth rate
A.2.1 The metabolic load of protein and the Hwa model
To model the effect of the metabolic load on cell growth, we will expand on a model used by Hwa and co-workers [14]. For conciseness, we will call the original model the Hwa model.
Consider a cell where the total number of proteins in the cell is $N$. The synthesis of these proteins requires two sets of processes: (i) the metabolic processes responsible for synthesizing the precursors (i.e. amino acids, etc) and (ii) the translation process. Proteins involved in the metabolic processes be referred to as the $P$ sector and number $N_{P}$. The proteins involved in the translational process will be referred to as the $R$ sector and number $N_{R}$. In addition to the P and R sectors, there is a third Q sector with protein number $N_{Q}$. The total protein number per cell is therefore:
$N=N_{R}+N_{P}+N_{Q}.$ (20)
The proteome fractions are defined $\Phi_{X}\equiv N_{X}/N$ and have the normalization condition:
$1=\Phi_{R}+\Phi_{P}+\Phi_{Q}.$ (21)
The key assumption in the Hwa model is that the abundances of the R and P sectors can change in size to accommodate changes in the nutrient quality and translation load associated with a particular growth condition [14]. In contrast, the Q sector has a fixed proteome fraction, $\Phi_{Q}$, irrespective of growth conditions. In the Hwa model, the size of these adjustable R and P sectors are chosen to optimize the growth rate $k$.
The condition for balanced growth requires that the overall protein output of the translation process match the growth rate:
$kN=k_{R}N_{*R},$ (22)
where $k_{R}$ is the effective translation rate per protein and $N_{*R}$ is the number of productive R sector proteins, which is subset of the total number $N_{R}$:
$N_{R}=N_{*R}+N_{0R},$ (23)
and $N_{0R}$ represents unproductive R sector protein. We can rewrite Eq. 22 in terms of the proteome fraction:
$k=k_{R}(\Phi_{R}-\Phi_{0R}).$ (24)
In the expression above, we will assume that the parameters $k_{R}$ and $\Phi_{0R}$ are fixed, but the total fraction $\Phi_{R}$ is chosen to optimize the growth rate $k$.
For any productive sectors $i$, we will write analogous equations to Eq. 24 linking sector fraction size $\Phi_{i}$ to function:
$k=k_{i}(\Phi_{i}-\Phi_{0i}),$ (25)
where, as before, $\Phi_{0i}$ represents a fixed-size fraction of unproductive protein.
To determine the unknown optimum growth rate and sector sizes, Eq. 25 can be re-written:
$\Phi_{*i}=\Phi_{i}-\Phi_{0i}=\textstyle\frac{k}{k_{i}},$ (26)
and then summed over all sectors (excluding Q):
$\Phi_{*}=k\sum_{i\neq Q}k_{i}^{-1},$ (27)
where we define the total productive fraction of the proteome:
$\Phi_{*}\equiv\sum_{i\neq Q}\Phi_{*i}=1-\Phi_{0},$ (28)
and $\Phi_{0}$ represents the total fraction of unproductive protein:
$\Phi_{0}=\Phi_{Q}+\sum_{i\neq Q}\Phi_{0i},$ (29)
including the entire Q sector. Since $\Phi_{Q}$ and the $\Phi_{0i}$ are all assumed to be fixed, Eq. 27 determines the growth rate. (In E. coli, Hwa and coworkers estimate that $\Phi_{*}\approx 0.55$.)
To understand the meaning of Eq. 27, we first define an ideal growth rate as
$k_{\rm ideal}^{-1}=\sum_{i\neq Q}k_{i}^{-1},$ (30)
which would be the growth rate in the absence of unproductive protein; however, due to the presence of the unproductive protein, the growth rate is proportional to the productive fraction:
$k=k_{\rm ideal}\Phi_{*}.$ (31)
The optimal protein fractions can be determined using Eq. 26:
$\Phi_{i}=\textstyle\frac{k_{\rm ideal}}{k_{i}}\Phi_{*}+\Phi_{0i}.$ (32)
How does the growth rate change when the unproductive protein fraction is changed by $\delta\Phi$? The productive fraction is reduced:
$\Phi_{*}\rightarrow\Phi_{*}^{\prime}=\Phi_{*}-\delta\Phi.$ (33)
The ratio of the new growth rate $k^{\prime}$ to the original is therefore:
$\frac{k^{\prime}}{k}=1-\textstyle\frac{\delta\Phi}{\Phi_{*}.}$ (34)
Note that our generalized Hwa model is written for arbitrary number of functional sectors $i\neq Q$ and the key determinant of the change in growth rate is the fraction of productive protein $\Phi_{*}$. This equation will be used to model the fitness cost of the metabolic load.
A.2.2 The metabolic load of mRNA
What is the cost of transcription? It is perhaps useful to first consider the estimates of biosynthetic cost of macromolecules in the cell in descending order in E. coli [38]:
Macromolecule Biosynthetic cost
($10^{9}$ ATP)
Protein 4.5
Phospholidid 3.2
RNA 1.6
Lipopolysacchride 3.8
DNA 0.35
Peptidoglycan 0.17
Glycogen 0.03
So clearly the cost of RNA is itself not insignificant. Although a significant fraction of the RNA is rRNA rather than mRNA, the mRNA itself in E. coli undergoes multiple rounds of transcription due to its short lifetime, increasing its cost to what is required to synthesize the molecules observed in the E. coli cell at any time $t$. Furthermore, transcription is dependent on protein enzymes, which themselves must be synthesized. We therefore conclude from this estimate that the cost of transcription is likely a significant determinant of the metabolic load.
Experimentally, Kafri and coworkers have measured the fitness cost of transcription and translation independently using the DAmP (Decreased Abundance by mRNA Perturbation) system in Saccharomyces cerevisiae [39]. As expected, they report that the metabolic cost of transcription is comparable to translation and that the reduction in growth rate is linear in transcription, in close analogy to Eq. 34.
A.2.3 Metabolic load in the RLTO model
To produce a minimal model to study the trade-off between robustness and metabolic load, we must consider both the metabolic cost of transcription and translation. We will write that the metabolic load (in protein equivalents) associated with gene $i$ is:
$\delta N_{i}=\lambda\mu_{m,i}+\mu_{p,i},$ (35)
@@ A.2.3 Metabolic load in the RLTO model
where $\lambda$ is the message cost, the metabolic load associated with an mRNA molecule relative to a single protein molecule of the gene product. $\mu_{m,i}$ is the mean number of messages transcribed per cell cycle (mRNA molecules per cell cycle) for gene $i$. $\mu_{p,i}$ is the mean number of protein translated per cell cycle for gene $i$. We will describe the mean protein number in terms of the translation efficiency $\varepsilon_{i}$, the number of proteins translated per message:
$\mu_{p,i}=\varepsilon_{i}\mu_{m,i}.$ (36)
How does the cell growth rate change due to the metabolic load associated with the expression of gene $i$? The change in the metabolic load is:
$\delta\Phi_{i}=\textstyle\frac{\delta N_{i}}{N},$ (37)
where $N$ represents the total metabolic load of all components of the cell, in units of protein equivalents. Using Eq. 34, the resulting change in growth rate is:
$\frac{k}{k_{0}}=1-\textstyle\frac{(\lambda+\varepsilon_{i})\mu_{m,i}}{\Phi_{*}N},$ (38)
where $k_{0}$ is the growth rate in the absence of the metabolic load of gene $i$.
In our analysis, the exact size of the total metabolic load $N$ will not be important. In the interest of simplicity we will therefore adsorb the productive fraction $\Phi_{*}$ into an effective total metabolic load:
$N_{0}\equiv\Phi_{*}N,$ (39)
and write a concise relation between the load from gene $i$ and the growth rate:
$\frac{k}{k_{0}}=1-\textstyle\frac{(\lambda+\varepsilon_{i})\mu_{m,i}}{N_{0}}.$ (40)
This equation has an intuitive interpretation: growth slows in proportion to the relative added metabolic load. Since $\Phi_{*}$ is order unity, we will ignore the distinction between the $N$ and $N_{0}$ quantities hence forth.
@@
Although the global parameters $N_{0}$ and $\lambda$ provide an intuitive representation of the model, the relative growth rate depends on fewer parameters. Let $k$ and $k_{0}$ be the growth rates in the presence and absence of the metabolic load of gene $i$. The relative growth rate is:
$\frac{k}{k_{0}}=1-(\Lambda+E_{i})\mu_{m,i},$ (41)
@@
where we have introduced two new reduced parameters: the relative load, defined as $\Lambda\equiv\lambda/N_{0}$, represents the ratio of the metabolic load of a single message to the total load and the relative translation efficiency, defined $E_{i}\equiv\varepsilon_{i}/N_{0}$, which is the ratio of the number of proteins translated per message to the total metabolic load $N_{0}$. (Note that due to the high multiplicity $N_{0}\gg(\lambda+\varepsilon_{i})\mu_{m,i}$, we can ignore the distinction between $N_{0}$ and $N_{0}^{\prime}$ in the denominator.) If we neglect the difference between the total metabolic load and the number of proteins, the proteome fraction for gene $i$ is:
$\Phi_{i}=E_{i}\mu_{m,i}.$ (42)
@@
Both reduced parameters, $\Lambda$ and $E_{i}$ are extremely small. In E. coli, we estimate that both $\Lambda$ and $E_{i}$ are roughly $10^{-5}$ and they are smaller still for eukaryotic cells. (See Sec. A.18.)
A.2.4 Growth rate with stochastic arrest
@@
is the exponential mean, where $T$ is the stochastic cell cycle duration and $\mathbb{E}$ is the expectation operator [40, 29].
Let $P_{+}$ be the probability of growth. When the cells are growing, the cell cycle duration $\tau$ is determined by the metabolic load predictions (Eq. 40). The probability mass function is therefore:
$p_{T}(t)=\begin{cases}P_{+},&t=\tau\\ (1-P_{+}),&t\rightarrow\infty\end{cases}.$ (45)
Evaluating the expectation in Eq. 44 gives:
$\overline{T}=-k\ln P_{+}+\tau.$ (46)
Using Eq. 43, we can solve for the growth rate $k$:
$k=\tau^{-1}\ln(2P_{+}).$ (47)
As expected, the growth rate goes down as the probability of growth $P_{+}$ decreases, stopping completely at $P_{+}=\frac{1}{2}$. We can then compute the ratio of the growth with ($k$) and without arrest ($k_{0}$):
$\frac{k}{k_{0}}=1+\textstyle\frac{1}{\ln 2}\ln P_{+}.$ (48)
where $k_{0}$ is computed by evaluating Eq. 47 at $P_{+}=1.$
Figure S2: Four perspectives on the fitness landscape. In each landscape, the red circle represents the fitness optimum. The red dotted line represents the mean protein number equal to the protein threshold $n_{p}=10^{2}$. Here, fitness is quantified by the log growth rate: $s=\ln k_{\rm max}/k$. Panel A: Mean protein number versus noise. Panel B: Mean protein number versus translation efficiency. Panel C: Mean message versus protein numbers. Panel D: Message number versus translation efficiency.
A.2.5 RLTO growth rate
@@
$\ln P_{+}$ $=$ $\sum_{i\in{\cal E}}\ln\left(1-{\rm Pr}\{N_{p,i}<n_{p,i}\}\right),$ (50)
$\approx$ $-\sum_{i\in{\cal E}}{\rm Pr}\{N_{p,i}<n_{p,i}\},$ (51)
$=$ $-\sum_{i\in{\cal E}}\textstyle\gamma(\frac{\mu_{m,i}}{\ln 2},\frac{n_{p,i}}{\varepsilon_{i}\ln 2}),$ (52)
where $\gamma$ is the regularized incomplete gamma function and the CDF of the gamma distribution (see  A.1.2 and A.1.4).
A.2.6 When the ln approximation is avoided…
The approximation discussed in the previous section is extremely well justified at the optimal central dogma parameters; however, there are a set of figures where we cannot use it. In the fitness landscape figures (Fig. S2), we compute the fitness not just at the optimal values but far from them. When cell arrest has a large effect on the growth rate, we cannot approximate the natural log with a series expansion, and we must use the full expression in Eq. 50.
A.2.7 Single-gene equation
By summing the fitness losses from the metabolic load and cell arrest (Eqs. 40, 48, and 52), we can write an expression for the growth rate including contributions from essential gene $i$:
$\textstyle\frac{k}{k_{0}}$ $=$ $1-\textstyle\frac{1}{N_{0}}(\lambda+\varepsilon_{i})\mu_{m,i}+...$ (53)
$-\textstyle\frac{1}{\ln 2}\gamma(\frac{\mu_{m,i}}{\ln 2},\frac{n_{p,i}}{\varepsilon\ln 2}),$
where the second term on the RHS represents the fitness loss due to the metabolic load and the third term represents the fitness loss due to stochastic cell arrest due to protein $i$ falling below threshold. The fitness landscape for different gene expression parameters is shown from four different perspectives in Fig. S2. From this point forward, we will drop the subscript $i$ for the sake of brevity unless otherwise noted.
A.2.8 Summary of RLTO parameter values for figures
The parameter values for the RLTO model used for each figure in the paper are shown in Tab. S1.